# Thousands of AI Authors on the Future of AI

### v1

@@ 4.3 How likely is AI to cause human extinction?
Statistics |
2022 Result |
2023 Result |
What probability do you put on future AI advances causing human extinction or similarly permanent and severe disempowerment of the human species? |
N; Mean (SD); Median (IQR) |
149; 15.7% (22.1%); 5% (19%) |
1321; 16.2% (23%); 5% (19%) |
What probability do you put on human inability to control future advanced AI systems causing human extinction or similarly permanent and severe disempowerment of the human species? |
N; Mean (SD); Median (IQR) |
162; 20.5% (26.2%); 10% (29%) |
661; 19.4% (26%); 10% (29%) |
What probability do you put on future AI advances causing human extinction or similarly permanent and severe disempowerment of the human species within the next 100 years?444Bold font was not present in the original questions; it has been added here to emphasize the differences between wordings. Any individual respondent did not see more than one of these questions. |
N; Mean (SD); Median (IQR) |
Not asked |
655; 14.4% (22.2%); 5% (19.9%) |
@@ 5.2.2 Participation
As with any survey, our results could be skewed by participation bias, if participants had systematically different opinions than those who chose not to participate. We sought to minimize this effect by aiming to maximize response rate, and by limiting cues about the survey content available before opting to take the survey. We looked for evidence of significant response bias at the survey level and question level, and did not find any. (see Appendix D for more detail)
@@ A.2.1 Amount of thinking about the issue
“Very little” or “a little” thought (n = 621) |
“A lot” or “a great deal” of thought (n = 1113) |
Percentage of these respondents who thought that the probability of an extremely good outcome was greater than 10% |
68% |
41% |
Percentage of these respondents who thought that the probability of an extremely bad outcome was greater than 10% |
34% |
71% |
@@ Appendix C Full Descriptions of Milestones
Short name |
Full description from question |
Physically install wiring in a house |
@@
Given a one-sentence description of the task and given the same information you would give a human to perform this task (such as information about the house), physically install the electrical wiring in a new home, without more input from humans. |
Finetune LLM |
@@
Given a one-sentence description of the task, download and finetune an existing open source LLM, without more input from humans. The fine-tune must improve the performance of the LLM on some predetermined benchmark metric. |
Find and patch security flaw |
@@
Given a one-sentence description of the task and no more input from humans, find and patch a security flaw in an open source project with over 100,000 users. |
Run ML study and write paper |
@@
Given a one sentence description of a research question in machine learning, conduct a study that would inform the answer to that question and write a paper of a quality that could be accepted at a leading machine learning conference, without more input from humans. |
Replicate ML paper |
@@
Given a study published at a leading machine learning conference, replicate the study without more input from humans. The replication must meet the standards of the ML Reproducibility Challenge (https://paperswithcode.com/rc2022). |
Build payment processing website |
@@
Given a set of specifications, build a website from scratch that can handle payment processing, including the frontend, backend, and secure payment integration, without more input from humans. |
Long unsolved math problem |
@@
Given a list of long-standing unsolved problems in mathematics, such as the Millennium Prize problems or one the problems in “Unsolved Problems on Mathematics for the 21st Century,” solve one without more input from humans. |
Translate text from a newly discovered language using a bilingual document |
@@
Translate a text written in a newly discovered language into English as well as a team of human experts, using a single other document in both languages (like a Rosetta stone). Suppose all of the words in the text can be found in the translated document, and that the language is a difficult one. |
Translate new language speech using films with subtitles and existing bilingual data |
@@
Translate speech in a new language given only unlimited films with subtitles in the new language. Suppose the system has access to training data for other languages, of the kind used now (e.g. same text in two languages for many languages and films with subtitles in many languages). |
Translate text nearly as well as a bilingual but untrained translator |
@@
Perform translation about as good as a human who is fluent in both languages but unskilled at translation, for most types of text, and for most popular languages (including languages that are known to be difficult, like Czech, Chinese and Arabic). |
Offer phone banking services, including unique tasks, on par with human operators |
@@
Provide phone banking services as well as human operators can, without annoying customers more than humans. This includes many one-off tasks, such as helping to order a replacement bank card or clarifying how to use part of the bank website to a customer. |
Classify unseen objects into categories after training on different but similar classes |
@@
Correctly group images of previously unseen objects into classes, after training on a similar labeled dataset containing completely different classes. The classes should be similar to the ImageNet classes. |
Recognize a new object in varied settings after seeing it just once |
@@
One-shot learning: see only one labeled image of a new object, and then be able to recognize the object in real world scenes, to the extent that a typical human can (i.e. including in a wide variety of settings). For example, see only one image of a platypus, and then be able to recognize platypuses in nature photos. The system may train on labeled images of other objects. Currently, deep networks often need hundreds of examples in classification tasks1, but there has been work on one-shot learning for both classification2and generative tasks.3 1: Lake et al. (2015). Building Machines That Learn and Think Like People 2: Koch (2015). Siamese Neural Networks for One-Shot Image Recognition 3: Rezende et al. (2016). One-Shot Generalization in Deep Generative Models |
Create a 3D model and realistic video from a new angle of a scene |
@@
See a short video of a scene, and then be able to construct a 3D model of the scene good enough to create a realistic video of the same scene from a substantially different angle. For example, constructing a short video of walking through a house from a video taking a very different path through the house. |
Transcribe speech in noise and varied accents on par with humans |
@@
Transcribe human speech with a variety of accents in a noisy environment as well as a typical human can. |
Output written text as a recording indistinguishable from a voice actor |
@@
Take a written passage and output a recording that can’t be distinguished from a voice actor, by an expert listener. |
Prove and generate math theorems publishable in top journals |
@@
Routinely and autonomously prove mathematical theorems that are publishable in top mathematics journals today, including generating the theorems to prove. |
Win Putnam Math Competition (problems with known but very difficult answers) |
@@
Perform as well as the best human entrants in the Putnam competition—a math contest whose questions have known solutions, but which are difficult for the best young mathematicians. |
Beat humans at Go (after same # games of training) |
@@
Defeat the best Go players, training only on as many games as the best Go players have played. For reference, DeepMind’s AlphaGo has probably played a hundred million games of self-play, while Lee Sedol has probably played 50,000 games in his life.1 1: Lake et al. (2015). Building Machines That Learn and Think Like People |
Beat ( 50% games) best Starcraft 2 players given only video of screen |
@@
Beat the best human Starcraft 2 players at least 50% of the time, given a video of the screen. Starcraft 2 is a real time strategy game characterized by:
Match human novice in any new computer game in <10 min |
@@
Play a randomly selected computer game, including difficult ones, about as well as a human novice, after playing the game less than 10 minutes of game time. The system may train on other games. |
Outperform humans in new Angry Birds levels |
@@
Play new levels of Angry Birds better than the best human players. Angry Birds is a game where players try to efficiently destroy 2D block towers with a catapult. For context, this is the goal of the IJCAI Angry Birds AI competition.1 1: aibirds.org |
Beat pro testers in all Atari games |
@@
Outperform professional game testers on all Atari games using no game-specific knowledge. This includes games like Frostbite, which require planning to achieve sub-goals and initially posed problems for deep Q-networks.1, 2 1: Mnih et al. (2015). Human-level control through deep reinforcement learning 2: Lake et al. (2015). Building Machines That Learn and Think Like People |
Beat novices in 50% of Atari games after 20 min play |
@@
Outperform human novices on 50% of Atari games after only 20 minutes of training play time and no game specific knowledge. For context, the original Atari playing deep Q-network outperforms professional game testers on 47% of games,1 but used hundreds of hours of play to train.2 1: Mnih et al. (2015). Human-level control through deep reinforcement learning 2: Lake et al. (2015). Building Machines That Learn and Think Like People |
@@
Fold laundry as well and as fast as the median human clothing store employee |
Fold laundry as well and as fast as the median human clothing store employee. |
Beat fastest human runners in a 5km city streets race using bipedal robot body |
@@
Beat the fastest human runners in a 5 kilometer race through city streets using a bipedal robot body. |
Build any LEGO set using using non-specialized robotics, with instructions |
@@
Physically assemble any LEGO set given the pieces and instructions, using non-specialized robotics hardware. For context, Fu 20161 successfully joins single large LEGO pieces using model based reinforcement learning and online adaptation. 1: Fu et al. (2016). One-Shot Learning of Manipulation Skills with Online Dynamics Adaptation and Neural Network Priors |
Efficiently sort large number lists beyond training size |
@@
Learn to efficiently sort lists of numbers much larger than in any training set used, the way Neural GPUs can do for addition,1 but without being given the form of the solution. For context, the original Neural Turing Machines could not do this,2 but Neural Programmer-Interpreters3 have been able to do this by training on stack traces (which contain a lot of information about the form of the solution). 1: Kaiser & Sutskever (2015). Neural GPUs Learn Algorithms 2: Zaremba & Sutskever (2015). Reinforcement Learning Neural Turing Machines 3: Reed & de Freitas (2015). Neural Programmer-Interpreters |
Write readable Python code for algorithms like quicksort from specs and examples |
@@
Write concise, efficient, human-readable Python code to implement simple algorithms like quicksort. That is, the system should write code that sorts a list, rather than just being able to sort lists. Suppose the system is given only:
Answer Googleable factoid questions better than expert (w/ web) |
@@
Answer any “easily Googleable” factoid questions posed in natural language better than an expert on the relevant topic (with internet access), having found the answers on the internet. Examples of factoid questions:
Answer Googleable but open-ended factual questions better than expert (w/ web) |
@@
Answer any “easily Googleable” factual but open ended question posed in natural language better than an expert on the relevant topic (with internet access), having found the answers on the internet. Examples of open ended questions:
Answer undecided factual questions |
@@
Give good answers in natural language to factual questions posed in natural language for which there are no definite correct answers. For example: “What causes the demographic transition?”, “Is the thylacine extinct?”, “How safe is seeing a chiropractor?” |
Write high-grade, unique high school history essays without plagiarizing |
@@
Write an essay for a high-school history class that would receive high grades and pass plagiarism detectors. For example answer a question like “How did the whaling industry affect the industrial revolution?” |
Create songs that can hit the US Top 40 (full audio file) |
@@
Compose a song that is good enough to reach the US Top 40. The system should output the complete song as an audio file. |
Fake new song indistinguishable from a specific artist’s work by expert listeners |
@@
Produce a song that is indistinguishable from a new song by a particular artist, e.g. a song that experienced listeners can’t distinguish from a new song by Taylor Swift. |
Write a novel or story that could land on the NYT best-seller list |
@@
Write a novel or short story good enough to make it to the New York Times best-seller list. |
Concisely and completely explain AI’s computer game moves |
@@
For any computer game that can be played well by a machine, explain the machine’s choice of moves in a way that feels concise and complete to a layman. |
Play poker well enough to win the World Series of Poker |
Play poker well enough to win the World Series of Poker. |
Deduce and symbolize physical laws (e.g. Newtonian mechanics) of a virtual world |
@@
After spending time in a virtual world, output the differential equations governing that world in symbolic form. For example, the agent is placed in a game engine where Newtonian mechanics holds exactly and the agent is then able to conduct experiments with a ball and output Newton’s laws of motion. |

### v2

@@ 4.3 How likely is AI to cause human extinction?
Statistics | 2022 Result | 2023 Result | |
What probability do you put on future AI advances causing human extinction or similarly permanent and severe disempowerment of the human species? | N; Mean (SD); Median (IQR) | 149; 15.7% (22.1%); 5% (19%) | 1321; 16.2% (23%); 5% (19%) |
What probability do you put on human inability to control future advanced AI systems causing human extinction or similarly permanent and severe disempowerment of the human species? | N; Mean (SD); Median (IQR) | 162; 20.5% (26.2%); 10% (29%) | 661; 19.4% (26%); 10% (29%) |
What probability do you put on future AI advances causing human extinction or similarly permanent and severe disempowerment of the human species within the next 100 years?444Bold font was not present in the original questions; it has been added here to emphasize the differences between wordings. Any individual respondent did not see more than one of these questions. | N; Mean (SD); Median (IQR) | Not asked | 655; 14.4% (22.2%); 5% (19.9%) |
@@ 5.2.2 Participation
As with any survey, our results could be skewed by participation bias, if participants had systematically different opinions than those who chose not to participate. We sought to minimize this effect by aiming to maximize response rate, and by limiting cues about the survey content available before opting to take the survey. We looked for evidence of response bias at the survey level and question level for some questions, and did not find any that would affect the results to a large extent (see Appendix D for more detail).
@@ A.2.1 Amount of thinking about the issue
“Very little” or “a little” thought (n = 621) | “A lot” or “a great deal” of thought (n = 1113) | |
Percentage of these respondents who thought that the probability of an extremely good outcome was greater than 10% | 68% | 41% |
Percentage of these respondents who thought that the probability of an extremely bad outcome was greater than 10% | 34% | 71% |
@@ Appendix C Full Descriptions of Milestones
Short name | Full description from question |
@@
Physically install wiring in a house | Given a one-sentence description of the task and given the same information you would give a human to perform this task (such as information about the house), physically install the electrical wiring in a new home, without more input from humans. |
@@
Finetune LLM | Given a one-sentence description of the task, download and finetune an existing open source LLM, without more input from humans. The fine-tune must improve the performance of the LLM on some predetermined benchmark metric. |
@@
Find and patch security flaw | Given a one-sentence description of the task and no more input from humans, find and patch a security flaw in an open source project with over 100,000 users. |
@@
Run ML study and write paper | Given a one sentence description of a research question in machine learning, conduct a study that would inform the answer to that question and write a paper of a quality that could be accepted at a leading machine learning conference, without more input from humans. |
@@
Replicate ML paper | Given a study published at a leading machine learning conference, replicate the study without more input from humans. The replication must meet the standards of the ML Reproducibility Challenge (https://paperswithcode.com/rc2022). |
@@
Build payment processing website | Given a set of specifications, build a website from scratch that can handle payment processing, including the frontend, backend, and secure payment integration, without more input from humans. |
@@
Long unsolved math problem | Given a list of long-standing unsolved problems in mathematics, such as the Millennium Prize problems or one the problems in “Unsolved Problems on Mathematics for the 21st Century,” solve one without more input from humans. |
@@
Translate text from a newly discovered language using a bilingual document | Translate a text written in a newly discovered language into English as well as a team of human experts, using a single other document in both languages (like a Rosetta stone). Suppose all of the words in the text can be found in the translated document, and that the language is a difficult one. |
@@
Translate new language speech using films with subtitles and existing bilingual data | Translate speech in a new language given only unlimited films with subtitles in the new language. Suppose the system has access to training data for other languages, of the kind used now (e.g. same text in two languages for many languages and films with subtitles in many languages). |
@@
Translate text nearly as well as a bilingual but untrained translator | Perform translation about as good as a human who is fluent in both languages but unskilled at translation, for most types of text, and for most popular languages (including languages that are known to be difficult, like Czech, Chinese and Arabic). |
@@
Offer phone banking services, including unique tasks, on par with human operators | Provide phone banking services as well as human operators can, without annoying customers more than humans. This includes many one-off tasks, such as helping to order a replacement bank card or clarifying how to use part of the bank website to a customer. |
@@
Classify unseen objects into categories after training on different but similar classes | Correctly group images of previously unseen objects into classes, after training on a similar labeled dataset containing completely different classes. The classes should be similar to the ImageNet classes. |
@@
Recognize a new object in varied settings after seeing it just once | One-shot learning: see only one labeled image of a new object, and then be able to recognize the object in real world scenes, to the extent that a typical human can (i.e. including in a wide variety of settings). For example, see only one image of a platypus, and then be able to recognize platypuses in nature photos. The system may train on labeled images of other objects. Currently, deep networks often need hundreds of examples in classification tasks1, but there has been work on one-shot learning for both classification2and generative tasks.3 1: Lake et al. (2015). Building Machines That Learn and Think Like People 2: Koch (2015). Siamese Neural Networks for One-Shot Image Recognition 3: Rezende et al. (2016). One-Shot Generalization in Deep Generative Models |
@@
Create a 3D model and realistic video from a new angle of a scene | See a short video of a scene, and then be able to construct a 3D model of the scene good enough to create a realistic video of the same scene from a substantially different angle. For example, constructing a short video of walking through a house from a video taking a very different path through the house. |
@@
Transcribe speech in noise and varied accents on par with humans | Transcribe human speech with a variety of accents in a noisy environment as well as a typical human can. |
@@
Output written text as a recording indistinguishable from a voice actor | Take a written passage and output a recording that can’t be distinguished from a voice actor, by an expert listener. |
@@
Prove and generate math theorems publishable in top journals | Routinely and autonomously prove mathematical theorems that are publishable in top mathematics journals today, including generating the theorems to prove. |
@@
Win Putnam Math Competition (problems with known but very difficult answers) | Perform as well as the best human entrants in the Putnam competition—a math contest whose questions have known solutions, but which are difficult for the best young mathematicians. |
@@
Beat humans at Go (after same # games of training) | Defeat the best Go players, training only on as many games as the best Go players have played. For reference, DeepMind’s AlphaGo has probably played a hundred million games of self-play, while Lee Sedol has probably played 50,000 games in his life.1 1: Lake et al. (2015). Building Machines That Learn and Think Like People |
@@
Beat ( 50% games) best Starcraft 2 players given only video of screen | Beat the best human Starcraft 2 players at least 50% of the time, given a video of the screen. Starcraft 2 is a real time strategy game characterized by: • Continuous time play • Huge action space • Partial observability of enemies • Long term strategic play, e.g. preparing for and then hiding surprise attacks. |
@@
Match human novice in any new computer game in <10 min | Play a randomly selected computer game, including difficult ones, about as well as a human novice, after playing the game less than 10 minutes of game time. The system may train on other games. |
@@
Outperform humans in new Angry Birds levels | Play new levels of Angry Birds better than the best human players. Angry Birds is a game where players try to efficiently destroy 2D block towers with a catapult. For context, this is the goal of the IJCAI Angry Birds AI competition.1 1: aibirds.org |
@@
Beat pro testers in all Atari games | Outperform professional game testers on all Atari games using no game-specific knowledge. This includes games like Frostbite, which require planning to achieve sub-goals and initially posed problems for deep Q-networks.1, 2 1: Mnih et al. (2015). Human-level control through deep reinforcement learning 2: Lake et al. (2015). Building Machines That Learn and Think Like People |
@@
Beat novices in 50% of Atari games after 20 min play | Outperform human novices on 50% of Atari games after only 20 minutes of training play time and no game specific knowledge. For context, the original Atari playing deep Q-network outperforms professional game testers on 47% of games,1 but used hundreds of hours of play to train.2 1: Mnih et al. (2015). Human-level control through deep reinforcement learning 2: Lake et al. (2015). Building Machines That Learn and Think Like People |
@@
Fold laundry as well and as fast as the median human clothing store employee | Fold laundry as well and as fast as the median human clothing store employee. |
@@
Beat fastest human runners in a 5km city streets race using bipedal robot body | Beat the fastest human runners in a 5 kilometer race through city streets using a bipedal robot body. |
@@
Build any LEGO set using using non-specialized robotics, with instructions | Physically assemble any LEGO set given the pieces and instructions, using non-specialized robotics hardware. For context, Fu 20161 successfully joins single large LEGO pieces using model based reinforcement learning and online adaptation. 1: Fu et al. (2016). One-Shot Learning of Manipulation Skills with Online Dynamics Adaptation and Neural Network Priors |
@@
Efficiently sort large number lists beyond training size | Learn to efficiently sort lists of numbers much larger than in any training set used, the way Neural GPUs can do for addition,1 but without being given the form of the solution. For context, the original Neural Turing Machines could not do this,2 but Neural Programmer-Interpreters3 have been able to do this by training on stack traces (which contain a lot of information about the form of the solution). 1: Kaiser & Sutskever (2015). Neural GPUs Learn Algorithms 2: Zaremba & Sutskever (2015). Reinforcement Learning Neural Turing Machines 3: Reed & de Freitas (2015). Neural Programmer-Interpreters |
@@
Write readable Python code for algorithms like quicksort from specs and examples | Write concise, efficient, human-readable Python code to implement simple algorithms like quicksort. That is, the system should write code that sorts a list, rather than just being able to sort lists. Suppose the system is given only: • A specification of what counts as a sorted list • Several examples of lists undergoing sorting by quicksort |
@@
Answer Googleable factoid questions better than expert (w/ web) | Answer any “easily Googleable” factoid questions posed in natural language better than an expert on the relevant topic (with internet access), having found the answers on the internet. Examples of factoid questions: • “What is the poisonous substance in Oleander plants?” • “How many species of lizard can be found in Great Britain?” |
@@
Answer Googleable but open-ended factual questions better than expert (w/ web) | Answer any “easily Googleable” factual but open ended question posed in natural language better than an expert on the relevant topic (with internet access), having found the answers on the internet. Examples of open ended questions: • “What does it mean if my lights dim when I turn on the microwave?” • “When does home insurance cover roof replacement?” |
@@
Answer undecided factual questions | Give good answers in natural language to factual questions posed in natural language for which there are no definite correct answers. For example: “What causes the demographic transition?”, “Is the thylacine extinct?”, “How safe is seeing a chiropractor?” |
@@
Write high-grade, unique high school history essays without plagiarizing | Write an essay for a high-school history class that would receive high grades and pass plagiarism detectors. For example answer a question like “How did the whaling industry affect the industrial revolution?” |
@@
Create songs that can hit the US Top 40 (full audio file) | Compose a song that is good enough to reach the US Top 40. The system should output the complete song as an audio file. |
@@
Fake new song indistinguishable from a specific artist’s work by expert listeners | Produce a song that is indistinguishable from a new song by a particular artist, e.g. a song that experienced listeners can’t distinguish from a new song by Taylor Swift. |
@@
Write a novel or story that could land on the NYT best-seller list | Write a novel or short story good enough to make it to the New York Times best-seller list. |
@@
Concisely and completely explain AI’s computer game moves | For any computer game that can be played well by a machine, explain the machine’s choice of moves in a way that feels concise and complete to a layman. |
Play poker well enough to win the World Series of Poker | Play poker well enough to win the World Series of Poker. |
@@
Deduce and symbolize physical laws (e.g. Newtonian mechanics) of a virtual world | After spending time in a virtual world, output the differential equations governing that world in symbolic form. For example, the agent is placed in a game engine where Newtonian mechanics holds exactly and the agent is then able to conduct experiments with a ball and output Newton’s laws of motion. |